Ab initio calculation of the photoabsorption cross section of NO_{2} by the two-dimensional reflection method.

A. DELON, B. KIRMSE, R. JOST, LCMI/CNRS BP166, 38042 Grenoble Cedex 9, France; \underline{J. Liévin}, Laboratoire de Chimie Physique Moléculaire, Université Libre de Bruxelles, CPI 160/09, B-1050 Bruxelles, BELGIUM.

The C_{2v} potential energy surfaces of the X^{2}A_{1}, A^{2}B_{2} and B^{2}B_{1} states of NO_{2} are calculated at a high level of theory (CASSCF/CMRCI/cc-pVTZ), together with the corresponding dipole moment surfaces for the transitions X −A and X −B. All surfaces are fitted to a quadratic polynomial form in the symmetric stretching and bending coordinates, around the equilibrium geometry of the ground state. These ab initio data are used to calculate the relative and absolute photoabsorption cross sections by applying the two-dimensional reflection method ^{a}, modified for taking the quadratic form of the potential energy surfaces into account. The theoretical cross section is found in good agreement with the experimental spectrum. The sensitivity of the calculated cross section with the level of ab initio theory will be discussed.

asee R. Schinke, ”Photodissociation dynamics”, Cambridge University Press (1993).